You have the following information about means and standard deviations of several investment options:

Investment Mean StDev
A 13 19
B 12 18
C 8 20

Your client wants to minimize the chance of having a less than 0% return. What is the safety first ratio of the optimal investment for the client?

Enter answer accurate to three decimal places.
Answer #1

Threshold mean = 0

Safety First (SF) ratio = (Expected return – Threshold return)/Standard deviation

For A, SF ratio = (13 – 0)/19 = 0.684

For B, SF ratio = (12 – 0)/18 = 0.667

For C, SF ratio = (8 – 0)/20 = 0.40

Portfolio A has the highest SF ratio and therefore should be selected.